# OpenOcc: Open Vocabulary 3D Scene Reconstruction via Occupancy Representation

### v1

@@ I Introduction
Reconstructing 3D scenes with high-level semantics from digital images and videos is a crucial area of research in 3D computer vision and robotics, which has been widely used in various fields such as visual navigation [1, 2], and robotic grasping [3, 4]. However, the intrinsic connections and underlying homologies between scene reconstruction and understanding have been greatly neglected. These two tasks are often treated as separate studies, where the bottleneck lies.
@@
However, it is not feasible to directly fuse these two types of work for the robotic structure perception, owing to two facts: 1) Current reconstruction methods [18, 20, 21] only perform better on 3D object surface reconstruction, which can not be extended to complex or large-scale indoor scenes. 2) Owing to the potential noise of language embedding, the open vocabulary segmentation predictions are inconsistent and ambiguous. Directly distilling these inconsistent features may lead to the degeneracy of the representation capabilities for language fields, which causes floaters in volume rendering.
To tackle this question, we propose OpenOcc, a novel framework unifying the scene reconstruction and open vocabulary understanding utilizing neural radiance field. To better learn the implicit representation, we construct the language, geometric, and appearance fields with separate multi-resolution feature gird. Instead of predicting a sign distance function (SDF), we utilize the occupancy representation surface as the decision boundary of a binary occupancy classifier. This leads to a significant reduction in memory requirements and improves the inference efficiency compared to the dense point-based and mesh-based representation. Moreover, the small objects and thin structures can be better reconstructed. To achieve open vocabulary scene understanding, we distill the 2D language embeddings into a 3D language field inspired by LeRF [16]. To address the issue of shape-radiance ambiguity, we have developed geometric constraints that incorporate depth information. Owing to the potential noise in the language embedding, the predictions from open vocabulary segmentation exhibit inconsistencies and ambiguities. Thus, we propose an SCP method to dynamically reallocate the weights of the corresponding language features in the language field updating, which can significantly improve the accuracy of 2D segmentation results.
@@ II Related Work
Neural implicit feature fields have emerged as a method to translate 2D features into a grounded 3D representation. CLIP-Fields [40] fuse CLIP [31] embeddings of crops into point clouds using a contrastively supervised field. OpenScene [41] and VL-Maps [2] build a 3D volume of language features aligned with the pre-trained language segmentation encoders [34, 35]. These methods are significantly constrained as they do not encode the intrinsic scene geometry, relying instead on an external point cloud for query operations. LERF [16] and VL-Fields [42] advocate the construction of a language field through volume rendering and the incorporation of the 2D language feature [31, 34] into the 3D implicit field, which is beneficial to obtain the dense and continuous scene perception results. However, owing to the potentially noisy nature of language embeddings, the predictions from open vocabulary segmentation exhibit inconsistencies and ambiguity. In this paper, we propose a semantic-aware confidence propagation method to dynamically reallocate the weights of the corresponding language features in the language field updating. The mistake observation of language embeddings can be efficiently filtered.
@@ III-A Occupancy Field Generation
To better reconstruct details of the 3D scene geometry structure, we utilize the multi-resolution feature grid via occupancy representation instead of a single scale inspired by Go-Surf [26]. We encode the scene geometry into a multi-level feature grid $\mathbb{V}_{\theta}=\{V^{l}\}$, where $l\in\{0,1,2,...,L\}$ encode the geometry details from coarse to fine. The coarse-level features can fastly initialize scene geometry presentation to fill the hole. Furthermore, the geometry residual of the high-frequency details will be refined by the fine-level features grid.
@@ III-B Multi-view Consistent Language Fields
$\small P\left(\tilde{s}_{t}\mid z_{1:t}\right)=\frac{\left(1-P\left(s_{t}\mid z_{t}\right)\right)P\left(z_{t}\right)\left(1-P\left(s_{t-1}\mid z_{1:t-1}\right)\right)}{(1-P(s_{t}))P\left(z_{t}\mid z_{1:t-1}\right)}.$ (10)
@@ III-C Training Loss
$\footnotesize\mathcal{L}(\hat{o}_{\mathbf{x}_{i}},o_{\mathbf{x}_{i}})=-[o_{\mathbf{x}_{i}}\cdot\log(\hat{o}_{\mathbf{x}_{i}})+(1-o_{\mathbf{x}_{i}})\cdot\log(1-\hat{o}_{\mathbf{x}_{i}})]$ (14)
@@ IV-A Experimental Setup
Datasets
We train and evaluate our model mainly on three popular public benchmarks: Replica [44], ScanNet [45], and Matterport3D [17]. These three datasets encompass the major situations of indoor scenes: the first two represent regular indoor scenes, and the last represents complex and large-scale indoor scenes, which all provide a series of posed RGB-D images. To more accurately and authentically showcase the segmentation capabilities of these methods on previously unaddressed classes, we also evaluate our method on the Scannet-200 dataset [46].
Metrics
The evaluation of the reconstructed 3D structure involves assessing six standard metrics as defined in [47]: Accuracy (Acc), Completeness (Comp), Chamfer-$\mathcal{L}_{1}$ (C-$\mathcal{L}_{1}$), Precision (Prec), Recall, and F-score. The F-score is computed using a threshold of 5cm. For scene understanding, we compute the two most common evaluation metrics, mean intersection-over-union (mIou) and mean accuracy (mAcc), which are used to evaluate the segmentation performance and accuracy of models. There is no unified and accurate evaluation metric for open vocabulary scene perception, so we mainly perform qualitative comparisons and discussions to demonstrate the zero-shot capability of our proposed method.
Implementation Details
We run our OpenOcc on a Tesla A6000 GPU for about 30 to 40 minutes(depending on the scene size) to finish the reconstruction and perception, which takes roughly 30GB of memory in total. For all our experiments, we sample 6144 rays per patch and sample 132 points along each ray. We further train our model for 10k iterations with a learning rate of $1\times 10^{-2}$ and $1\times 10^{-3}$ for MLP decoders and feature grids. Due to the universality of our approach, unconstrained by specific semantic segmentation methods, we choose pre-trained OpenSeg  [48] for 2D feature distillation in our experiment part.
@@ IV-B Experimental Results
We evaluate our OpenOcc on all classes of the Replica, Scannet-200 validation set and Matterport3D test set. In Tab. I and Tab. II, we compare our methods with the most closely related work on zero-shot methods. Due to the inherent shape-radiance ambiguity problem associated with the LeRF approach, it can not acquire accurate geometric structures on the validation dataset. Consequently, a comparison with it has not been conducted. Our model integrates 3D scene reconstruction with open vocabulary segmentation, exhibiting superior performance relative to OpenScene, particularly in the segmentation of small objects or long-tail objects, as evidenced in Tab. I. Since reconstructing a large number of scenes will consume too much resource and time, we establish a new validation subset by randomly extracting 10% of data(about 30 scenes) from the Scannet validation set, and a similar operation is performed on the Matterport3D test set. For fairness, we also evaluate OpenScene under the same subset split.
@@
TABLE III: Quantitative evaluation of reconstruction map. We evaluate the 3D reconstruction performance of our proposed method against previous works on ScanNet dataset with six represent metrics. “***” denotes the results generated from the source code with the default parameters.
@@ IV-D Memory and Time
We exhibit the merits of our occupancy-based reconstruction and representation method in terms of low storage requirements and high efficiency, shown in Fig. 1. Compared with the SDF-based NeRF method, our occupancy-based NeRF variant reduces training time by a significant 74.7%, from 33min down to just 8min. Moreover, implicit fields typically require 4.4GB of memory, whereas our explicit language feature occupancy grid takes only 1.1GB of memory, a reduction of 75%.

### v2

@@ I Introduction
Reconstructing 3D scenes with high-level semantics from digital images or videos is a crucial research area in 3D computer vision and robotics, which has been widely used in various fields such as visual navigation [1, 2], and robotic grasping [3, 4]. However, the intrinsic connections and underlying homologies between scene reconstruction and understanding have been neglected. These two tasks are often treated as separate studies, where the bottleneck lies.
@@
However, it is not feasible to directly fuse these two types of work for the robotic structure perception, owing to two facts: 1) Current reconstruction methods [18, 20, 21] only perform better on 3D object surface reconstruction, which can not be extended to complex or large-scale indoor scenes. 2) Owing to the potential noise of language embedding, the open vocabulary segmentation predictions are ambiguous. Directly distilling these inconsistent features may lead to the degeneracy of the representation capabilities for language fields, which causes floaters in volume rendering.
To tackle this question, we propose OpenOcc, a novel framework unifying the scene reconstruction and open vocabulary understanding utilizing NeRF. To better learn the implicit representation, we construct the language, geometric, and appearance fields with separate multi-resolution feature gird. Instead of predicting a sign distance function (SDF), we utilize the occupancy representation surface as the decision boundary of a binary occupancy classifier. This leads to a significant reduction in memory requirements and improves the inference efficiency compared to the dense point-based and mesh-based representation. Moreover, the small objects and thin structures can be better reconstructed. To achieve open vocabulary scene understanding, we distill the 2D language embeddings into a 3D language field inspired by LeRF [16]. To address the issue of shape-radiance ambiguity, we have developed geometric constraints that incorporate depth information. Owing to the potential noise in the language embedding, the predictions from open vocabulary segmentation exhibit inconsistencies and ambiguities. Thus, we propose an SCP method to dynamically reallocate the weights of the corresponding language features in the language field updating, which can significantly improve the accuracy of 2D segmentation results.
@@ II Related Work
Neural implicit feature fields have emerged as a method to translate 2D features into a grounded 3D representation. CLIP-Fields [40] fuse CLIP [31] embeddings of crops into point clouds using a contrastively supervised field. OpenScene [41] and VL-Maps [2] build a 3D volume of language features aligned with the pre-trained language segmentation encoders [34, 35]. These methods are significantly constrained as they do not encode the intrinsic scene geometry, relying instead on an external point cloud for query operations. LERF [16] and VL-Fields [42] advocate the construction of a language field through volume rendering and the incorporation of the 2D language feature [31, 34] into the 3D implicit field, which is beneficial to obtain the dense and continuous scene perception. However, owing to the potentially noisy nature of language embeddings, the predictions from open vocabulary segmentation exhibit inconsistencies and ambiguity. In this paper, we propose a semantic-aware confidence propagation method to dynamically reallocate the weights of the corresponding language features in the language field updating. The mistake observation of language embeddings can be efficiently filtered.
@@ III-A Occupancy Field Generation
To better reconstruct details of the 3D scene geometry structure, we utilize the multi-resolution feature grid [43] via occupancy representation instead of a single scale inspired by Go-Surf [26]. We encode the scene geometry into a multi-level feature grid $\mathbb{V}_{\theta}=\{V^{l}\}$, where $l\in\{0,1,2,...,L\}$ encode the geometry details from coarse to fine. The coarse-level features can fastly initialize scene geometry presentation to fill the hole. Furthermore, the geometry residual of the high-frequency details will be refined by the fine-level features grid.
@@ III-B Multi-view Consistent Language Fields
$P\left(\tilde{s}_{t}\mid z_{1:t}\right)=\frac{\left(1-P\left(s_{t}\mid z_{t}\right)\right)P\left(z_{t}\right)\left(1-P\left(s_{t-1}\mid z_{1:t-1}\right)\right)}{(1-P(s_{t}))P\left(z_{t}\mid z_{1:t-1}\right)}.$
@@ III-C Training Loss
$\mathcal{L}(\hat{o}_{\mathbf{x}_{i}},o_{\mathbf{x}_{i}})=-[o_{\mathbf{x}_{i}}\cdot\log(\hat{o}_{\mathbf{x}_{i}})+(1-o_{\mathbf{x}_{i}})\cdot\log(1-\hat{o}_{\mathbf{x}_{i}})]$ (14)
@@ IV-A Experimental Setup
Datasets. We train and evaluate our model mainly on three popular public benchmarks: Replica [45], ScanNet [46], and Matterport3D [17]. These three datasets encompass the major situations of indoor scenes: the first two represent regular indoor scenes, and the last represents complex and large-scale indoor scenes, which all provide a series of posed RGB-D images. To more accurately and authentically showcase the segmentation capabilities of these methods on previously unaddressed classes, we also evaluate our method on the Scannet-200 dataset [47].
Metrics. The evaluation of the reconstructed 3D structure involves assessing six standard metrics as defined in [48]: Accuracy (Acc), Completeness (Comp), Chamfer-$\mathcal{L}_{1}$ (C-$\mathcal{L}_{1}$), Precision (Prec), Recall, and F-score. The F-score is computed using a threshold of 5cm. For scene understanding, we compute the two most common evaluation metrics, mean intersection-over-union (mIou) and mean accuracy (mAcc), which are used to evaluate the segmentation performance and accuracy of models. There is no unified and accurate evaluation metric for open vocabulary scene perception, so we mainly perform qualitative comparisons and discussions to demonstrate the zero-shot capability of our proposed method.
Implementation Details. We run our OpenOcc on a Tesla A6000 GPU for about 30 to 40 minutes(depending on the scene size) to finish the reconstruction and perception, which takes roughly 30GB of memory in total. For all our experiments, we sample 6144 rays per patch and sample 132 points along each ray. We further train our model for 10k iterations with a learning rate of $1\times 10^{-2}$ and $1\times 10^{-3}$ for MLP decoders and feature grids. Due to the universality of our approach, unconstrained by specific semantic segmentation methods, we choose pre-trained OpenSeg  [49] for 2D feature distillation in our experiment part.
@@ IV-B Experimental Results
We evaluate our OpenOcc on all classes of the Replica, Scannet-200 validation set and Matterport3D test set. In Tab. I and Tab. II, we compare our methods with the most closely related work on zero-shot methods. Due to the original LeRF approach not utilizing depth supervision, it inherently suffers from shape-radiance ambiguity. Consequently, we have not compared our method with it in this part. Our model integrates 3D scene reconstruction with open vocabulary segmentation, exhibiting superior performance relative to OpenScene, particularly in the segmentation of small objects or long-tail objects, as evidenced in Tab. I. Since reconstructing a large number of scenes will consume too much resource and time, we establish a new validation subset by randomly extracting 10% of data(about 30 scenes) from the Scannet validation set, and a similar operation is performed on the Matterport3D test set. For fairness, we also evaluate OpenScene under the same subset split.
@@
TABLE III: Quantitative evaluation of reconstruction map. We evaluate the 3D reconstruction performance of our proposed method against previous works on ScanNet dataset with six represent metrics. “*” denotes the results generated from the source code with the default parameters.
@@ IV-D Memory and Time
We exhibit the merits of our occupancy-based reconstruction and representation method in terms of low storage requirements and high efficiency, shown in Fig. 1. The occupancy-based representation allows for the efficient exclusion of invalid sampled spaces. Compared with the SDF-based NeRF method, our occupancy-based NeRF variant reduces training time by a significant 74.7%, from 33min down to just 8min. Moreover, implicit fields typically require 4.4GB of memory, whereas our explicit language feature occupancy grid takes only 1.1GB of memory, a reduction of 75%.